OpenStudy (anonymous):

WILL GIVE MEDALS What is the equation of y = x3 with the given transformations? vertical stretch by a factor of 3, horizontal shift 4 units to the right, vertical shift 3 units down

OpenStudy (anonymous):

@Skiller8860 @hartnn

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

transformations of a function: \(\Large{ y=a(bx^3−c)+d}\) a is vertical stretch/compression a| > 1 stretches |a| < 1 compresses a < 0 flips the graph upside down b is horizontal stretch/compression |b| > 1 compresses |b| < 1 stretches b < 0 flips the graph left-right c is horizontal shift c < 0 shifts to the right c > 0 shifts to the left d is vertical shift d > 0 shifts upward d < 0 shifts downward (accdg to math is fun) apply your given transformations.

OpenStudy (anonymous):

@Pleasehelpasap12345 okay so which one?

OpenStudy (anonymous):

whAT IS THE QUESTION

OpenStudy (anonymous):

@Pleasehelpasap12345 C,A,A,A,D,D,B,A,C,A, 3x+3y=10 infinite solutions

OpenStudy (anonymous):

dON'T INDERSTAND

OpenStudy (anonymous):

sorry for caps

OpenStudy (anonymous):

@Precious56 this is the question What is the equation of y = x3 with the given transformations?

OpenStudy (anonymous):

do you know how to do graph

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay, so draw one and then on the x axis g up 3 on the y axis

OpenStudy (anonymous):

go up sorry

OpenStudy (anonymous):

lol ok

OpenStudy (anonymous):

hope you do good

OpenStudy (anonymous):

but what would be the equation?

OpenStudy (anonymous):

well if you do that it would be y=x and the numb you landed on

OpenStudy (anonymous):

hope that helps

OpenStudy (anonymous):

3f(x-4) - 3 = 3(x-4)^3 - 3

OpenStudy (anonymous):

@Pleasehelpasap12345 welcome

OpenStudy (anonymous):

can you help me

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

what is y=x-1wth an arrow 5=4-1

OpenStudy (anonymous):

it's linear equations

OpenStudy (anonymous):

hmmm one sec

OpenStudy (anonymous):

k

OpenStudy (anonymous):

x=5 y=4 @Precious56

OpenStudy (anonymous):

thanks I appreciate it emails me any time Saxqueen352@yahoo.com